# Sensor-fusion based Prognostics for Deep-space Habitats Exhibiting Multiple Unlabeled Failure Modes

### v3

@@ Abstract
Deep-space habitats are complex systems that must operate autonomously over extended durations without ground-based maintenance. These systems are vulnerable to multiple, often unknown, failure modes that affect different subsystems and sensors in mode-specific ways. Developing accurate remaining useful life (RUL) prognostics is challenging, especially when failure labels are unavailable and sensor relevance varies by failure mode. In this paper, we propose an unsupervised prognostics framework that jointly identifies latent failure modes and selects informative sensors using only unlabeled training data. The methodology consists of two phases. In the offline phase, we model system failure times using a mixture of Gaussian regressions and apply a novel Expectation-Maximization algorithm to cluster degradation trajectories and select mode-specific sensors. In the online phase, we extract low-dimensional features from the selected sensors to diagnose the active failure mode and predict RUL using a weighted regression model. We demonstrate the effectiveness of our approach on a simulated dataset that reflects deep-space telemetry characteristics and on a real-world engine degradation dataset, showing improved accuracy and interpretability over existing methods.
keywords:
Deep Space Habitats , Multiple Failure Modes , Remaining Useful Life , Sensor Fusion , Failure Diagnosis , Prognostics
journal: Acta Astronautica
[…]
@@ 1 Introduction
A deep-space habitat is a complex engineering system (CES), composed of multiple interacting subsystems that must work together to support human life and mission operations over long durations. These subsystems such as the environmental control and life support system, energy and power systems, thermal control, robotic agents, and structural components etc., must function reliably in harsh and isolated environments, where failure of any component can lead to mission-critical consequences [1, 2, 3, 4]. The complexity of such systems arises not only from the tight integration of these subsystems but also from their heterogeneous origins. Components are often developed by different manufacturers and tested independently, making it difficult to anticipate how they will behave once integrated. Additionally, the deep-space environment introduces operating conditions such as microgravity, radiation, and thermal extremes that differ significantly from terrestrial test settings. These novel conditions can give rise to failure modes that have never been observed during isolated ground-based validation. As a result, building a comprehensive library of failure modes becomes essential for predicting degradation behaviors (i.e., prognostics) during the mission. However, due to limited diagnostic access and the lack of labeled failure data, this library must often be inferred from unlabeled system observations.
@@
Figure 1: Example of a deep space habitat with multiple subsytems. Health of each subsystem is monitored by a subset of onboard sensors. The subset of informative sensors vary for different failure modes.
@@ 2 Literature Review
In reality, most CES can fail through multiple failure modes, reflecting the structure and function of their subsystems. Competing risk models have been widely used to represent such systems. Traditional approaches often rely on statistical models, such as the Cox proportional hazards model with Weibull baselines [29, 30]; a comprehensive review is provided in [31]. More recently, deep learning has also been used to model competing risks [32]. For example, both [33] and [34] both employ deep recurrent neural networks for recurrent event survival analysis with competing risks.
Competing risk models often produce survival or hazard rates that can be difficult to interpret in engineering contexts. As a result, many researchers have shifted toward methods that directly model sensor data to predict RUL, though few focus on multiple failure modes. It is common for researchers to develop general RUL prediction methodologies and then apply them to datasets with multiple failure modes. For example, [35] proposed a correlation and relative entropy feature engineering framework for complex systems and [36] proposed a deep convolutional neural network. In both cases, the researchers validated their methodologies on datasets with multiple failure modes. However, they do not consider multiple failure modes in an explicit manner. Failure to do so could result in highly biased predictions. Therefore, the development of RUL prediction methodologies that account for multiple failure modes is growing. [37] used a support vector machine for failure mode diagnosis with features extracted from a deep belief network, followed by RUL prediction using a particle filter. [38] trained distinct LSTM models for each failure mode. Once a physics-informed CNN classifier diagnoses the active failure mode, the appropriate LSTM model is utilized. Furthermore, [39] fit random coefficient models to failure mode-specific health indices. However, these articles do not incorporate a sensor selection algorithm and they rely on labeling of the failure modes determined either through visual inspection of domain knowledge. Consequently, researchers have begun addressing these gaps. [40] adopted a semi-supervised graph-based approach for feature extraction from partially labeled failures and used elastic net functional regression for sensor selection and RUL prediction. More recently, [41] proposed a time-series clustering algorithm to identify latent failure modes, followed by LSTM-based RUL prediction. Furthermore, [42] combined mixture of (log)-location scale regression with deep learning to predict RUL in a scenario with unknown failure modes. However, in both papers, sensor selection was limited to excluding sensors with obviously non-informative signals, such as constant or low-variance signals.
@@ 3 Overview of the Prognostics Framework
    Diagnosis: Prior to the diagnosis step, the active failure mode is unknown therefore, all selected sensors are utilized in this step. We apply multivariate functional principal component analysis (MFPCA) to signals from all selected sensors to extract a compact set of features, called MFPC scores, that represent the joint behavior of all sensors. We classify the active failure mode by finding its K-nearest neighbors and assigning the most common failure mode among them.
@@ 4 Offline Sensor Selection
We consider a scenario where the deep space habitat is monitored by $P$ different sensors. We assume there exists a historical dataset comprised of $P$-dimensional degradation-based sensor signals from monitoring $N$ similar samples. We denote $s_{i,p}(t)$ as the signal data from sensor $p\in\{1,...,P\}$ for system $i\in\{1,...,N\}$ observed at time $t\in[0,T]$, where $T=min(\{TTF_{i}\}_{i=1}^{N})$. By truncating all signals to the minimum failure time, we ensure that all signals in the training set have the same time domain and that they are all included when modeling. For this framework, we assume that the total number of failure modes, $K$ is known, but that the actual failure modes in the training dataset are unknown. This poses a challenge for CA-FPCA, which requires observed covariates for computation. To determine these covariates, we perform a sensor-wise clustering of the training samples. For each sensor, we perform FPCA for feature extraction and then use K-means clustering with $K$ clusters to label the signals from that sensor. Once the sensor-wise clustering is complete, the next step involves feature extraction by applying CA-FPCA and then regressing these features against the time-to-failure to identify the most informative subset of sensors for each failure mode. The next two subsections discuss details about the feature extraction (CA-FPCA), and the optimization (MGR-ASGL) model.
@@ 4.1 Feature Extraction: CA-FPCA
$s_{i,p}(t)=v_{i,p}^{k}(t)+\epsilon_{i,p}(t),i\in\mathcal{I}_{k}^{p}\vspace*{-1em}$ (1)
@@
$s_{i,p}(t)=\mu_{p}^{k}(t)+\sum_{m=1}^{\infty}\xi_{i,m,p}^{k}\phi_{m,p}^{k}(t)+\epsilon_{i,p}(t)\vspace*{-1em}$ (2)
@@ 4.2 Optimization of the MGR-ASGL model
Let $Y_{i}=\frac{\ln{TTF}_{i}-(1/N)\sum_{i=1}^{N}\ln{TTF}_{i}}{(1/(N-1))\sum_{i=1}^{N}\big{(}\ln{TTF}_{i}-(1/N)\sum_{i=1}^{N}\ln{TTF}_{i}\big{)}^{2}}$, for $i=1,2…,N$ denote the random variable corresponding to the standardized natural logarithm of the time-to-failure (i.e., $\ln{TTF}$) of system $i$. Furthermore, let $\boldsymbol{Z}_{i}$ denote a categorical random vector of size $K$ used to encode the unknown failure mode of system $i$. If failure mode $k$ is responsible for the failure of system $i$, which occurs with probability $\pi_{k}$, then $\boldsymbol{Z}_{i}[k]=1$ and the remaining elements are zero. In our model, we posit that the relationship between $Y_{i}$ and the observed features $x_{i}$ depends on which failure mode caused the failure of system $i$. Thus, we assume that the conditional distribution of $Y_{i}$ given $\boldsymbol{Z}_{i}[k]=1$ is Gaussian with PDF:
$f_{Y_{i}|\boldsymbol{Z}_{i}}(y_{i}|\boldsymbol{Z}_{i}[k]=1)=\frac{1}{\sqrt{2\pi}\sigma_{k}}\exp\Big{(}-\frac{1}{2\sigma_{k}^{2}}(y_{i}-\beta_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\beta}_{p,k})^{2}\Big{)}$ (3)
@@
$f_{Y_{i}}(y_{i})=\sum_{k=1}^{K}\pi_{k}\frac{1}{\sqrt{2\pi}\sigma_{k}}\exp\Big{(}-\frac{1}{2\sigma_{k}^{2}}(y_{i}-\beta_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\beta}_{p,k})^{2}\Big{)}$ (4)
@@
$-\ell(\Theta|\boldsymbol{Y})=-\sum_{i=1}^{N}\ln\Big{[}\sum_{k=1}^{K}\pi_{k}\rho_{k}\frac{1}{\sqrt{2\pi}}\exp\Big{(}-\frac{1}{2}(y_{i}\rho_{k}-\varphi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,k})^{2}\Big{)}\Big{]}$ (5)
@@
$\begin{split}-\ell_{C}(\Theta|\boldsymbol{Y},\mathbb{Z})&=-\ln\prod_{i=1}^{N}\prod_{k=1}^{K}f_{Y_{i}}(y_{i})^{\boldsymbol{Z}_{i}[k]}=-\sum_{i=1}^{N}\sum_{k=1}^{K}\boldsymbol{Z}_{i}[k]\ln\Big{[}\pi_{k}\rho_{k}\frac{1}{\sqrt{2\pi}}\\ &\times\exp\Big{(}-\frac{1}{2}(y_{i}\rho_{k}-\varphi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,k})^{2}\Big{)}\Big{]}\end{split}$ (6)
@@
    E-step: Given $\Theta^{(m)}$, the expectation of $-\ell_{C}(\Theta|\boldsymbol{Y},\mathbb{Z})$ is given as:
    $\begin{split}Q(\Theta,\Theta^{(m)})&=\mathbb{E}[-\ell_{C}(\Theta|\boldsymbol{Y},\mathbb{Z})|\boldsymbol{Y},\Theta^{(m)}]=-\sum_{i=1}^{N}\sum_{k=1}^{K}\gamma_{i,k}\ln\Big{[}\pi_{k}\rho_{k}\frac{1}{\sqrt{2\pi}}\\ &\times\exp\Big{(}-\frac{1}{2}(y_{i}\rho_{k}-\varphi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,k})^{2}\Big{)}\Big{]}\end{split}$ (8)
@@
    $\begin{split}\gamma_{i,k}&=\mathbb{E}(\boldsymbol{Z}_{i}[k]|Y_{i}=y_{i},\Theta^{(m)})=\Pr(\boldsymbol{Z}_{i}[k]=1|Y_{i}=y_{i},\Theta^{(m)})\\ &=\frac{\pi_{k}^{(m)}\rho_{k}^{(m)}\frac{1}{\sqrt{2\pi}}\exp\Big{(}-\frac{1}{2}(y_{i}\rho_{k}^{(m)}-\varphi_{0,k}^{(m)}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,k}^{(m)})^{2}\Big{)}}{\sum_{l=1}^{K}\pi_{l}^{(m)}\rho_{l}^{(m)}\frac{1}{\sqrt{2\pi}}\exp\Big{(}-\frac{1}{2}(y_{i}\rho_{l}^{(m)}-\varphi_{0,l}^{(m)}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,l}^{(m)})^{2}\Big{)}}\end{split}$ (9)
@@
    $\min_{\Theta}\Big{\{}Q(\Theta,\Theta^{(m)})+\lambda\sum_{k=1}^{K}\pi_{k}\Big{(}\alpha\sum_{p=1}^{P}\|\boldsymbol{\varphi}_{p,k}\|_{1}+(1-\alpha)\sum_{p=1}^{P}\sqrt{q_{p}}\|\boldsymbol{\varphi}_{p,k}\|_{2}\Big{)}\Big{\}}$ (10)
@@
    • (a)
      Step 1: Improvement with respect to $\boldsymbol{\pi}$
@@
    • (b)
      Step 2: Improvement with respect to $\boldsymbol{\rho}$ and $\boldsymbol{\varphi}$
@@
      $\begin{split}\min_{\rho_{k},\varphi_{0,k},\boldsymbol{\varphi}_{1,k},\dots,\boldsymbol{\varphi}_{P,k}}&\Big{\{}-\sum_{i=1}^{N}\gamma_{i,k}\big{[}\ln\pi_{k}^{(m+1)}+\ln\rho_{k}-\frac{1}{2}(y_{i}\rho_{k}-\varphi_{0,k}\\ &-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,k})^{2}\big{]}\ +\lambda\pi_{k}^{(m+1)}\Big{(}\alpha\sum_{p=1}^{P}\|\boldsymbol{\varphi}_{p,k}\|_{1}\\ &+(1-\alpha)\sum_{p=1}^{P}\sqrt{q_{p}}\|\boldsymbol{\varphi}_{p,k}\|_{2}\Big{)}\Big{\}}\end{split}$ (12)

### v4

@@ Abstract
Deep-space habitats are complex systems that must operate autonomously over extended durations without ground-based maintenance. These systems are vulnerable to multiple, often unknown, failure modes that affect different subsystems and sensors in mode-specific ways. Developing accurate remaining useful life (RUL) prognostics is challenging, especially when failure labels are unavailable and sensor relevance varies by failure mode. In this paper, we propose an unsupervised prognostics framework that jointly identifies latent failure modes and selects informative sensors using only unlabeled training data. The methodology consists of two phases. In the offline phase, we model system failure times using a mixture of Gaussian regressions and apply an Expectation-Maximization algorithm to cluster degradation trajectories and select mode-specific sensors. In the online phase, we extract low-dimensional features from the selected sensors to diagnose the active failure mode and predict RUL using a weighted regression model. We demonstrate the effectiveness of our approach on a simulated dataset that reflects deep-space telemetry characteristics and on a real-world engine degradation dataset, showing improved accuracy and interpretability over existing methods.
{IEEEkeywords}
Deep-space Habitats, Failure Diagnosis, Multiple Failure Modes, Prognostics, Remaining Useful Life, Sensor Fusion
@@ 1 INTRODUCTION
Deep-space habitat is a complex engineering system (CES), composed of multiple interacting subsystems that must work together to support human life and mission operations over long durations. These subsystems such as the environmental control and life support system, energy and power systems, thermal control, robotic agents, and structural components etc., must function reliably in harsh and isolated environments, where failure of any component can lead to mission-critical consequences [1, 2]. The complexity of such systems arises not only from the tight integration of these subsystems but also from their heterogeneous origins. Components are often developed by different manufacturers and tested independently, making it difficult to anticipate how they will behave once integrated. Additionally, the deep-space environment introduces operating conditions such as microgravity, radiation, and thermal extremes that differ significantly from terrestrial test settings. These novel conditions can give rise to failure modes that have never been observed during isolated ground-based validation. As a result, building a comprehensive library of failure modes becomes essential for predicting degradation behaviors (i.e., prognostics) during the mission. However, due to limited diagnostic access and the lack of labeled failure data, this library must often be inferred from unlabeled system observations.
@@
Figure 1: Example of a deep-space habitat with multiple subsytems. Health of each subsystem is monitored by a subset of onboard sensors. The subset of informative sensors vary for different failure modes.
@@ 2 LITERATURE REVIEW
In reality, most CES experience multiple failure modes, reflecting the structure and function of their subsystems. Competing risk models have been widely used to represent such systems. Traditional approaches often rely on statistical models, such as the Cox proportional hazards model with Weibull baselines [27, 28]; a comprehensive review is provided in [29]. More recently, deep learning has also been used to model competing risks [30]. For example, both [31] and [32] both employ deep recurrent neural networks for recurrent event survival analysis with competing risks.
Competing risk models often produce survival or hazard rates that can be difficult to interpret in engineering contexts. As a result, many researchers have shifted toward methods that directly model sensor data to predict RUL, though few focus on multiple failure modes. Often, researchers develop general RUL prediction methodologies and then apply them to datasets with multiple failure modes. For example, [33] proposed a correlation and relative entropy feature engineering framework for complex systems and [34] proposed a deep convolutional neural network. In both cases, the researchers validated their methodologies on datasets with multiple failure modes. However, they do not consider multiple failure modes in an explicit manner. Failure to do so could result in highly biased predictions. Therefore, the development of RUL prediction methodologies that account for multiple failure modes is growing. [35] used a support vector machine for failure mode diagnosis with features extracted from a deep belief network, followed by RUL prediction using a particle filter. [36] trained distinct LSTM models for each failure mode. Once a physics-informed CNN classifier diagnoses the active failure mode, the appropriate LSTM model is utilized. Furthermore, [37] fit random coefficient models to failure mode-specific health indices. However, these articles do not incorporate a sensor selection algorithm and they rely on labeling of the failure modes determined either through visual inspection or domain knowledge. Consequently, researchers have begun addressing these gaps. [38] adopted a semi-supervised graph-based approach for feature extraction from partially labeled failures and used elastic net functional regression for sensor selection and RUL prediction. More recently, [39] proposed a time-series clustering algorithm to identify latent failure modes, followed by LSTM-based RUL prediction. Furthermore, [40] combined mixture of (log)-location scale regression with deep learning to predict RUL in a scenario with unknown failure modes. However, in both papers, sensor selection was limited to excluding sensors with obviously non-informative signals, such as constant or low-variance signals.
@@ 3 OVERVIEW OF THE PROGNOSTICS FRAMEWORK
    Diagnosis: Prior to the diagnosis step, the active failure mode is unknown. Therefore, all selected sensors are utilized in this step. First we apply multivariate functional principal component analysis (MFPCA) to signals from all selected sensors to extract a compact set of features, called MFPC scores, that represent the joint behavior of all sensors. Then we classify the active failure mode by finding its K-nearest neighbors and assigning the most common failure mode among them.
@@ 4 OFFLINE SENSOR SELECTION
We consider a scenario where the deep-space habitat is monitored by $P$ different sensors. We assume there exists a historical dataset comprised of $P$-dimensional degradation-based sensor signals from monitoring $N$ similar systems. We denote $s_{i,p}(t)$ as the signal data from sensor $p\in\{1,...,P\}$ for system $i\in\{1,...,N\}$ observed at time $t\in[0,T]$, where $T=min(\{TTF_{i}\}_{i=1}^{N})$. By truncating all signals to the minimum failure time, we ensure that all signals in the training set have the same time domain and that they are all included when modeling. For this framework, we assume that the total number of failure modes, $K$ is known, but that the actual failure labels in the training dataset are unknown. This poses a challenge for CA-FPCA, which requires observed covariates for computation. To determine these covariates, we perform a sensor-wise clustering of the training samples. For each sensor, we perform FPCA for feature extraction and then use K-means clustering with $K$ clusters to label the signals from that sensor. Once the sensor-wise clustering is complete, the next step involves feature extraction by applying CA-FPCA and then regressing these features against the time-to-failure to identify the most informative subset of sensors for each failure mode. The next two subsections discuss details about the feature extraction (CA-FPCA), and the optimization (MGR-ASGL) model.
@@ 4.1 Feature Extraction: CA-FPCA
$s_{i,p}(t)=v_{i,p}^{k}(t)+\epsilon_{i,p}(t),i\in\mathcal{I}_{k}^{p}\vskip-10.00002pt$ (1)
@@
$s_{i,p}(t)=\mu_{p}^{k}(t)+\sum_{m=1}^{\infty}\xi_{i,m,p}^{k}\phi_{m,p}^{k}(t)+\epsilon_{i,p}(t)\vskip-10.00002pt$ (2)
@@ 4.2 Optimization of the MGR-ASGL model
Let $Y_{i}=\frac{\ln{TTF}_{i}-(1/N)\sum_{i=1}^{N}\ln{TTF}_{i}}{(1/(N-1))\sum_{i=1}^{N}\big(\ln{TTF}_{i}-(1/N)\sum_{i=1}^{N}\ln{TTF}_{i}\big)^{2}}$, for i=1,2,Ni=1,2…,N denote the random variable corresponding to the standardized natural logarithm of the time-to-failure (i.e., $\ln{TTF}$) of system $i$. Furthermore, let $\boldsymbol{Z}_{i}$ denote a categorical random vector of size $K$ used to encode the unknown failure mode of system $i$. If failure mode $k$ is responsible for the failure of system $i$, which occurs with probability $\pi_{k}$, then $\boldsymbol{Z}_{i}[k]=1$ and the remaining elements are zero. In our model, we posit that the relationship between $Y_{i}$ and the observed features $x_{i}$ depends on which failure mode caused the failure of system $i$. Thus, we assume that the conditional distribution of $Y_{i}$ given $\boldsymbol{Z}_{i}[k]=1$ is Gaussian with PDF:
$\begin{split}f_{Y_{i}|\boldsymbol{Z}_{i}}(y_{i}|\boldsymbol{Z}_{i}[k]=1)=\frac{1}{\sqrt{2\pi}\sigma_{k}}&\exp\Big(-\frac{1}{2\sigma_{k}^{2}}(y_{i}-\beta_{0,k}\\ &-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\beta}_{p,k})^{2}\Big)\end{split}$ (3)
@@
$\begin{split}f_{Y_{i}}(y_{i})=\sum_{k=1}^{K}\pi_{k}\frac{1}{\sqrt{2\pi}\sigma_{k}}&\exp\Big(-\frac{1}{2\sigma_{k}^{2}}(y_{i}-\beta_{0,k}\\ &-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\beta}_{p,k})^{2}\Big)\end{split}$ (4)
@@
$\begin{split}-\ell(\Theta|\boldsymbol{Y})=-\sum_{i=1}^{N}\ln\Big[\sum_{k=1}^{K}\pi_{k}&\rho_{k}\frac{1}{\sqrt{2\pi}}\exp\Big(-\frac{1}{2}(y_{i}\rho_{k}\\ &-\varphi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,k})^{2}\Big)\Big]\end{split}$ (5)
@@
$\begin{split}-\ell_{C}(\Theta|\boldsymbol{Y},\mathbb{Z})&=-\ln\bigg(\prod_{i=1}^{N}\prod_{k=1}^{K}f_{Y_{i}}(y_{i})^{\boldsymbol{Z}_{i}[k]}\bigg)\\ &=-\sum_{i=1}^{N}\sum_{k=1}^{K}\boldsymbol{Z}_{i}[k]\ln\Big[\pi_{k}\rho_{k}\frac{1}{\sqrt{2\pi}}\exp\Big(\\ &-\frac{1}{2}(y_{i}\rho_{k}-\varphi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,k})^{2}\Big)\Big]\end{split}$ (6)
@@
    E-step: Given $\Theta^{(m)}$, the expectation of $\ell_{C}(\Theta|\boldsymbol{Y},\mathbb{Z})$ is given as:
    $\begin{split}Q(\Theta,\Theta^{(m)})&=\mathbb{E}[-\ell_{C}(\Theta|\boldsymbol{Y},\mathbb{Z})|\boldsymbol{Y},\Theta^{(m)}]\\ &=-\sum_{i=1}^{N}\sum_{k=1}^{K}\gamma_{i,k}\ln\Big[\pi_{k}\rho_{k}\frac{1}{\sqrt{2\pi}}\exp\Big(\\ &-\frac{1}{2}(y_{i}\rho_{k}-\varphi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,k})^{2}\Big)\Big]\end{split}$ (8)
@@
    $\begin{split}\gamma_{i,k}&:=\mathbb{E}\!\left(\boldsymbol{Z}_{i}[k]\mid Y_{i}=y_{i},\Theta^{(m)}\right)\\ &=\Pr\!\left(\boldsymbol{Z}_{i}[k]=1\mid Y_{i}=y_{i},\Theta^{(m)}\right)\\ &=\footnotesize\frac{\begin{aligned} &\pi_{k}^{(m)}\rho_{k}^{(m)}\frac{1}{\sqrt{2\pi}}\exp\!\Big(-\frac{1}{2}\big(y_{i}\rho_{k}^{(m)}-\varphi_{0,k}^{(m)}\\ &\hskip 34.14322pt-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,k}^{(m)}\big)^{2}\Big)\end{aligned}}{\begin{aligned} &\sum_{l=1}^{K}\pi_{l}^{(m)}\rho_{l}^{(m)}\frac{1}{\sqrt{2\pi}}\exp\!\Big(-\frac{1}{2}\big(y_{i}\rho_{l}^{(m)}-\varphi_{0,l}^{(m)}\\ &\hskip 34.14322pt-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,l}^{(m)}\big)^{2}\Big)\end{aligned}}\end{split}$ (9)
@@
    $\begin{split}\min_{\Theta}\Big\{Q(\Theta,\Theta^{(m)})&+\lambda\sum_{k=1}^{K}\pi_{k}\Big(\alpha\sum_{p=1}^{P}\|\boldsymbol{\varphi}_{p,k}\|_{1}\\ &+(1-\alpha)\sum_{p=1}^{P}\sqrt{q_{p}}\|\boldsymbol{\varphi}_{p,k}\|_{2}\Big)\Big\}\end{split}$ (10)
@@
    • Step 1: Improvement w.r.t. $\boldsymbol{\pi}$
@@
    • Step 2: Improvement w.r.t. $\boldsymbol{\rho}$ and $\boldsymbol{\varphi}$
@@
      $\begin{split}\min\Big\{&-\sum_{i=1}^{N}\gamma_{i,k}\big[\ln\pi_{k}^{(m+1)}+\ln\rho_{k}\\ &-\frac{1}{2}(y_{i}\rho_{k}-\varphi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\varphi}_{p,k})^{2}\big]\ \\ &+\lambda\pi_{k}^{(m+1)}\Big(\alpha\sum_{p=1}^{P}\|\boldsymbol{\varphi}_{p,k}\|_{1}\\ &+(1-\alpha)\sum_{p=1}^{P}\sqrt{q_{p}}\|\boldsymbol{\varphi}_{p,k}\|_{2}\Big)\Big\}\end{split}$ (12)